Circular No. 8195

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 R1 (LINEAR) An apparently asteroidal object reported by LINEAR (discovery ob-

servation below), and posted on the NEO Confirmation Page, has been reported to have cometary appearance on CCD images obtained by J. Tichá and M. Tichý (Kleť, 1.06-m KLENOT telescope; slightly diffuse object with a 6′′ coma on Sept. 5.08 UT, and asymmetric coma to the northwest on Sept. 6.05); and by J. E. McGaha (Tucson, AZ, 0.30-m f/10.0 Schmidt-Cassegrain reflector; 3′′ coma with a fan-shaped tail 8′′ long in p.a. 320^{o} on Sept. 6.39).

2003 UT  α_{2000}  δ_{2000}  Mag.
Sept.2.36756  3^{h}26^{m}08.^{s}86  −1^{o}31′55′′.5  18.3

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2003-R35.

T = 2003 July 9.968 TT  ω = 306.^{o}497
Ω = 356.282 }2000.0
q = 2.15420 AU  i = 149.675

POSSIBLE NOVA IN M33 M. Ganeshalingam and W. Li, University of California at Berkeley

report the discovery via LOTOSS (cf. IAUC 8192) of an apparent nova on unfiltered KAIT images on Sept. 1.4 (mag ∼ 16.9), 4.4 (mag ∼ 17.0), and 6.4 UT (mag ∼ 17.1). The nova is located at α = 1^{h}33^{m}51^{s}.16, δ = +30^{o}39′50′′.2 (equinox 2000.0), which is 3′′.3 east and 14′′.4 north of the nucleus of M33 (= NGC 598). A KAIT image taken on Aug. 24.4 showed nothing at this position (limiting mag ∼ 19.5).

CORRIGENDA On IAUC 8172, SUPERNOVA 2003gt IN NGC 6930, line 5, for 4′′.7

south read 9′′.7 south

On IAUC 8188, ϵ INDI B, final sentence, the numbers are simple brightness ratios, as a fraction of the brightness of ϵ Ind B, not as a percentage of the brightness of ϵ Ind B.

2003 September 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green